# Orca: Browsing at Scale Through User-Driven and AI-Facilitated Orchestration Across Malleable Webpages

### v1

@@ 8.1. Blessings and Curses of Spatial Information Interfaces
Finally, an overview across web content, or information, does not have to be a spatial one. Future research could investigate alternative structures, such as a linear grid (similar to a spreadsheet), where users can place webpages or other content in fixed cells, to understand the tradeoffs compared to spatial layouts.
@@ 8.2. Malleable User Interfaces for the Web
With rapidly advancing LLM capabilities and emerging agentic infrastructures, such as Model Context Protocol (MCP) and NLWeb, we anticipate AI agents will soon be able to navigate, operate, interpret, and extract information from most websites with greater reliability and depth (llm, 2024; Anthropic and contributors, 2025; Microsoft, 2025). To this end, we envision treating webpages not merely as visual interfaces, but as “services” or APIs that provide rich data, content, and functionality.
LLM-powered browsers can then serve as interface “composers” and leverage their generative capabilities to create custom, task-specific interfaces layered over the web-served content. Agentic pipelines serve as a transformation layer that connects the custom UI and data of varied formats and schemas (Anthropic and contributors, 2025; Microsoft, 2025). This would enable the creation of malleable, personalized, and evolving workspaces that can selectively incorporate information aligned with users’ tasks across “data sources” (Cao et al., 2025; Min et al., 2023).
8.3. Parallel and Collaborative Real-Time Multi-Agent Systems
Orca introduces more granular, contextualized, and multi-threaded web automation by allowing multiple agents to operate in parallel across pages to complete tasks.
However, feedback from participants revealed a need for better monitoring and control of real-time multi-agent systems. P6 suggested a global progress bar to track agents in one place, rather than scattered across the Web Canvas. P2 proposed breaking down and visualizing the agents’ detailed steps using an interactive diagram to make it easier for users to track and adjust their behavior (Jiang et al., 2023a; Kazemitabaar et al., 2024).
Therefore, future research can explore new mechanisms to monitor agent progress and enable timely user intervention. This could involve methods such as direct manipulation, demonstrations, or even a “group chat” interface for users to give intermediate feedback to one or more agents simultaneously (Lau, 2009).
@@ 9. Conclusion
We introduce Orca, a prototype web browser that reconceptualizes webpages as malleable materials and supports user-driven, AI-facilitated orchestration across them in the malleable browser space. With a spatial canvas interface, Orca provides features to enable browsing at scale across webpages, addressing the distributed nature of web-based informational tasks. Our evaluation shows that Orca reduces the effort required to manage complex browsing workflows while preserving user agency, increasing engagement, and stimulating users’ desire for information foraging. These findings highlight a promising direction for future browser designs that tightly integrate AI capabilities into the personal, user-driven informational tasks.
Appendix A Constructing Feedforward Options Lists from Parallel LLM Responses
@@ Appendix B HTML Distillation for Web Automation and Page Extraction
The id is a stable, three-character unique identifier derived from the element’s xpath, ensuring consistent reference across distillations.
@@
[ACTIVE] $>$ [\$m2l\$:click (Send email)] Send $<$
@@
That is, a button titled “Send” will be extracted, and clicking it will propagate the click event to the underlying element with id m2l.

### v2

@@ 7.3.2. Custom use of the spatial layout helps with diverse activities of the exploration process (RQ2)
7.3.5. Continuous visibility and batch controls help users identify and recover from AI failures (RQ5)
Because current LLM-based agents are stochastic and error-prone, failures are inevitable: they may misinterpret instructions, act on the wrong elements, or navigate to unintended pages (Huq et al., 2025).
To monitor and prune agent actions, P1 routinely range-selected webpages to be updated after each request and focused the camera on them (Figure 3.1), using the canvas as a visual checklist. They also appreciated that the camera view automatically snapped to newly created webpages, making new content and activity immediately visible. When Contextual Expansion produced pages that did not match expectations, P1 selected and deleted them at once. Similarly, P7 closely monitored agents and, when an agent tried to add an unwanted grocery item, deleted that agent, zoomed in on the relevant page, and completed the step manually instead.
Meanwhile, once a page moved outside the current viewport, participants had difficulty tracking updates or agent behaviors on that page. As P7 zoomed into a single page to take over actions with a focused view, they were surprised upon zooming back out to find that many other agents had already finished their tasks elsewhere on the canvas. Although they did not observe obvious errors, this surfaced a limitation of “off-canvas” agents operating on pages outside the user’s view. A future version could pause such off-canvas agents by default, especially for potentially dangerous actions; or add off-screen activity indicators on the minimap or screen edges, similar to cues used in digital maps.
Participants also adapted their strategies when a feature repeatedly failed. For example, when P4 wanted to see flights from adjacent days, they first tried Contextual Expansion with query “find flights +-2 days” and expected the four expanded pages to land on the correct dates. However, the model repeatedly attempted to open target pages by adding URL parameters—disallowed by Google Flights. As a remedy, P4 used global undo (Command + Z) to revert the operation across pages, then duplicated the original page, range-selected the copies, and commanded “navigate to days +-2 days”. Orca then correctly deployed one agent per page, each navigating to a different, expected day.
@@ 8.1. Blessings and Curses of Spatial Information Interfaces
Thrid, an overview across web content, or information, does not have to be a spatial one. Future research could investigate alternative structures, such as a linear grid (similar to a spreadsheet), where users can place webpages or other content in fixed cells, to understand the tradeoffs compared to spatial layouts.
Finally, while Orca introduces a drastically different UI from today’s browsers, it is not intended to replace the tab-based paradigm. Instead, our exploration surfaces a complementary interface for user-driven information-foraging tasks in the era of AI. The Web Canvas layout can coexist with a traditional tabbed UI—for example, zooming into a page could transition to a tabbed view, whereas zooming out would return to a canvas where the page is surrounded by others. A future deployment study could examine how, given such dual modes and fluid transitions, users appropriate different views for different tasks and preferences.
@@ 8.2. Malleable User Interfaces for the Web
With rapidly advancing LLM capabilities and emerging agentic infrastructures, such as Model Context Protocol (MCP) and NLWeb, we anticipate AI agents will soon be able to navigate, operate, interpret, and extract information from most websites with greater reliability and depth (103; Anthropic and contributors (2025); Microsoft (2025); 66). To this end, we envision treating webpages not merely as visual interfaces, but as “services” or APIs that provide rich data, content, and functionality.
LLM-powered browsers can then serve as interface “composers” and leverage their generative capabilities to create custom, task-specific interfaces layered over the web-served content (Klokmose et al., 2015). Agentic pipelines serve as a transformation layer that connects the custom UI and data of varied formats and schemas (Anthropic and contributors, 2025; Microsoft, 2025). This would enable the creation of malleable, personalized, and evolving workspaces that can selectively incorporate information aligned with users’ tasks across “data sources” (Cao et al., 2025; Min et al., 2023).
8.3. Parallel and Browser-Level Multi-Agent Systems
Orcaintroduces granular, contextualized, multi-threaded web automation by allowing multiple agents to operate in parallel across pages to complete tasks. Because these browser-level agents can read and act on content across many tabs, they surface new privacy and security challenges, including what information they may access, how their actions are scoped across sites, and how cross-site data flows are governed at the browser level. In our current design, agents are invoked via explicit user commands on selected page subsets as their permitted context, and their actions remain visible and interruptible in real time, providing containment, oversight, and user control.
On the other hand, participant feedback revealed a need for better monitoring and control of real-time multi-agent systems. P6 suggested a global progress bar to track all agents, rather than scattered across the Web Canvas. P2 proposed breaking down and visualizing the agents’ detailed steps using an interactive diagram to make it easier for users to track and adjust their behavior (Jiang et al., 2023a; Kazemitabaar et al., 2024).
Therefore, future research can explore new mechanisms to monitor agent progress, manage permissions, and enable timely user intervention. This could involve methods such as direct manipulation, demonstrations, or even a “group chat” interface for users to give intermediate feedback to one or more agents simultaneously, complemented by workspace-level permission models and UI that explicitly scope which pages, data, and operations each agent is allowed to access (Lau, 2009). We see fine-grained visibility, controllability, and scoping as central to making browser-level multi-agent orchestration not only more usable, but also more privacy-preserving and secure, critical for it to scale and be adopted by millions.
@@ 9. Conclusion
We introduce Orca, a prototype web browser that reconceptualizes webpages as malleable materials and supports user-driven, AI-facilitated orchestration across them in the malleable browser space, aiding web-based informational work.
With a spatial canvas interface, Orca provides features to enable browsing at scale across webpages, addressing the distributed nature of web-based informational tasks. Our evaluation shows that Orca reduces the effort required to manage complex browsing workflows while preserving user agency, increasing engagement, and stimulating users’ desire for information foraging. These findings highlight a promising direction for future browser designs that tightly integrate AI capabilities into the personal, user-driven informational tasks.
Acknowledgements.
We thank members of the Foundation Interface Lab for their support and generous help during the production of the work.
Appendix A Constructing Feedforward Option Lists from Parallel LLM Responses
@@ Appendix B HTML Distillation for Web Automation and Page Extraction
The id is a stable, short, and unique identifier derived from the element’s xpath, ensuring consistent reference across distillations.
@@
[ACTIVE] >>> [\$m2l\$:click (Send email)] Send <<<
@@
That is, a button titled “Send” will be extracted, and clicking it will propagate the click event to the underlying element m2l.